OpenStudy (anonymous):

Find the amplitude, period, and phase shift of the function? y= 8sin(5/6x-pi/5)

OpenStudy (anonymous):

we know that the general formula is the following:\[y = A*\sin(\omega t +\theta)\]

OpenStudy (anonymous):

so in this case we can compute the following: - A = 8 -\[\theta\] (phase shift) = pi/5 - w = 5/6 (t)

OpenStudy (anonymous):

I forgot what w = ? to really, but if you have te formula for w, then you can computer t ^_^

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

compute*

OpenStudy (anonymous):

\(\omega = 2\pi f\) where \(f\) is the reciprocal of the period.

OpenStudy (anonymous):

thank you abhorsen ^_^

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

so in this case:\[T= 2\pi / w = 2\pi /5/6 = 7.5\] Correct me if I'm wrong

OpenStudy (anonymous):

T= 7.5 secs

OpenStudy (anonymous):

wait I forgot f lol

OpenStudy (anonymous):

nvm F ^^"

OpenStudy (anonymous):

Final Answer: - A = 8 - phase angle = pi/5 - T = 7.5 secs